# Semi-supervised Graph Anomaly Detection via Robust Homophily Learning

### v1

@@ Appendix C Description of Baselines
  • DOMINANT [6] applied the conventional autoencoder on a graph for GAD. It consists of an encoder layer and a decoder layer, which are designed to reconstruct the features and structure of the graph. The reconstruction errors from the features and the structural modules are combined as an anomaly score.
@@
  • AnomalyDAE [11] leverages the advanced autoencoder and an attribute autoencoder to learn both node embeddings and attribute embeddings jointly in a latent space by assigning the corresponding weight in the loss function. In addition, an attention mechanism is employed in the structure encoder to capture normal structural patterns more effectively.
@@
  • OCGNN [40] combines one-class SVM and GNNs, aiming at leveraging one-class classifiers and the powerful representation of GNNs. A one-class hypersphere learning objective is used to drive the training of the GNN. The samples that fall outside the hypersphere, meaning they deviate from the normal pattern, are defined as anomaly.
@@
  • AEGIS [5] designs a new graph neural layer to learn anomaly-aware node representations and further employ generative adversarial networks to detect anomalies among new data.The generator takes noise sampled from a prior distribution as input and aims to produce informative pseudo-anomalies. Meanwhile, the discriminator seeks to determine whether a given input is the representation of a normal node or a generated anomaly.
@@
  • GAAN [4] is based on a generative adversarial network where fake graph nodes are generated by a generator. To encode the nodes, they compute the sample covariance matrix for real nodes and fake nodes, and a discriminator is trained to recognize whether two connected nodes are from a real or fake node.
@@
  • TAM [34] adopts the local affinity-based approach as an anomaly measure. It learns tailored node representations for a new anomaly measure by maximizing the local affinity of nodes to their neighbors. TAM is optimized on truncated graphs where non-homophily edges are removed iteratively. The learned representations result in a significantly stronger local affinity for normal nodes than abnormal nodes.
@@
  • GGAD [36] generates pseudo anomaly nodes based on two important priors, including asymmetric local affinity and egocentric closeness, and trains a discriminative one-class classifier for semi-supervised GAD.
@@
  • GCN [16] obtains the node representations by aggregating information from neighbors, which can be seen as a special form of low-pass filter. The low-pass filter primarily preserves the similarity of node features during graph representation learning.
@@
  • BWGNN [39] reveals the ‘right shift’ phenomenon that the spectral energy distribution concentrates more on the node with high frequencies and less on the node with low frequencies, and proposes a band-pass filter to learn the effective node representation for supervised GAD.
@@ D.1 Homophily Distribution Results for the Remaining Datasets
D.4 Time Complexity Analysis
In this section, we analyze the time complexity of RHO. (1) The time complexity of the two-layer MLP used for initial feature transformation in RHO is $\mathcal{O}(Nd_{1}(M+d))$, where $N$ denotes the number of nodes, $M$ is the input feature dimension, and $d_{1}$ and $d$ represent the dimensions of the intermediate and output embeddings, respectively. (2) RHO employs two independent views, including the cross-channel and channel-wise views. Each view incurs a computational cost of $\mathcal{O}(|\mathcal{E}|d+Nd^{2})$, where $|\mathcal{E}|$ is the number of edges. The overall complexity of this component remains $\mathcal{O}(2*(|\mathcal{E}|d+Nd^{2}))$ in practice. (3) In GNA, the projection head has a complexity of $\mathcal{O}(Nd^{2})$, and the alignment loss computation has a complexity of $\mathcal{O}(Nbd)$, where $b$ denotes the batch size. Therefore, the overall time complexity is $\mathcal{O}((Nd_{1}(M+d))+2|\mathcal{E}|d+3Nd^{2}+Nbd))$.

### v2

@@ Appendix C Description of Baselines
  • DOMINANT [ding2019deep] applied the conventional autoencoder on a graph for GAD. It consists of an encoder layer and a decoder layer, which are designed to reconstruct the features and structure of the graph. The reconstruction errors from the features and the structural modules are combined as an anomaly score.
@@
  • AnomalyDAE [fan2020anomalydae] leverages the advanced autoencoder and an attribute autoencoder to learn both node embeddings and attribute embeddings jointly in a latent space by assigning the corresponding weight in the loss function. In addition, an attention mechanism is employed in the structure encoder to capture normal structural patterns more effectively.
@@
  • OCGNN [wang2021one] combines one-class SVM and GNNs, aiming at leveraging one-class classifiers and the powerful representation of GNNs. A one-class hypersphere learning objective is used to drive the training of the GNN. The samples that fall outside the hypersphere, meaning they deviate from the normal pattern, are defined as anomaly.
@@
  • AEGIS [ding2021inductive] designs a new graph neural layer to learn anomaly-aware node representations and further employ generative adversarial networks to detect anomalies among new data.The generator takes noise sampled from a prior distribution as input and aims to produce informative pseudo-anomalies. Meanwhile, the discriminator seeks to determine whether a given input is the representation of a normal node or a generated anomaly.
@@
  • GAAN [chen2020generative] is based on a generative adversarial network where fake graph nodes are generated by a generator. To encode the nodes, they compute the sample covariance matrix for real nodes and fake nodes, and a discriminator is trained to recognize whether two connected nodes are from a real or fake node.
@@
  • TAM [hezhe2023truncated] adopts the local affinity-based approach as an anomaly measure. It learns tailored node representations for a new anomaly measure by maximizing the local affinity of nodes to their neighbors. TAM is optimized on truncated graphs where non-homophily edges are removed iteratively. The learned representations result in a significantly stronger local affinity for normal nodes than abnormal nodes.
@@
  • GGAD [qiao2024generative] generates pseudo anomaly nodes based on two important priors, including asymmetric local affinity and egocentric closeness, and trains a discriminative one-class classifier for semi-supervised GAD.
@@
  • GCN [kipf2016semi] obtains the node representations by aggregating information from neighbors, which can be seen as a special form of low-pass filter. The low-pass filter primarily preserves the similarity of node features during graph representation learning.
@@
  • BWGNN [tang2022rethinking] reveals the ‘right shift’ phenomenon that the spectral energy distribution concentrates more on the node with high frequencies and less on the node with low frequencies, and proposes a band-pass filter to learn the effective node representation for supervised GAD.
  • PMP [zhuo2024partitioning] introduces a partitioning message passing scheme that separately aggregates information from homophilic and heterophilic neighbors using node-specific functions. This design enables the adaptive adjustment of information flow from different neighbor types, thereby effectively capturing structural heterogeneity and enhancing robustness to heterophily in graph-based fraud detection.
  • CONSISGAD [chen2024consistency] introduces a consistency-based framework for GAD under limited supervision. It leverages unlabeled data through a learnable data augmentation mechanism that injects controlled noise for consistency training. Additionally, by exploiting the variance in homophily distributions between normal and anomalous nodes, CONSISGAD employs a simplified GNN backbone to enhance discriminability and robustness against class imbalance.
@@ D.1 Homophily Distribution Results for the Remaining Datasets
D.4 Runtime Results
Table 4: Runtimes (in seconds) on the six datasets on CPU.
Methods Datasets
Reddit Photo Amazon Question T-Finance
DoMINANT 125 437 1592 740 10721
OCGNN 162 125 765 973 5717
AEGIS 166 417 1121 660 15258
TAM 432 165 4516 29280 17360
GGAD 368 136 1020 1125 9345
CONSISGAD 483 681 3259 483 20172
RHO 398 201 3358 675 5184
The running time, including both training and inference time, of RHO and six competing methods are shown in Table 4. Although our method can run on all datasets using an RTX 3090 GPU, some baseline methods may encounter out-of-memory (OOM) issues on the same GPU due to their different computational mechanisms, highlighting that RHO is more memory-efficient. To ensure a fair comparison, we report the runtime performance on the CPU for all methods. The runtime results demonstrate that RHO remains efficient even on the large-scale graph T-Finance, primarily due to the use of mini-batch processing in the GNA module. By adopting a mini-batch strategy, the complexity of GNA reduces to $O(b^{2}d)$ for a batch size of $b$, and the total per-epoch complexity becomes $O(Nbd)$ when averaged across all batches, which significantly reduces memory usage and computational overhead, thereby improving the scalability. In contrast, methods such as DOMINANT, TAM, and GGAD involve reconstruction or affinity calculations across all nodes, which typically incur a time complexity of $O(N^{2})$, causing significant computational overhead on large-scale graphs.